# A Tutorial on MIMO-OFDM ISAC: From Far-Field to Near-Field

### v1

@@ IV Far-field MIMO-OFDM ISAC
In this section, we elaborate on how to sense far-field targets based on the signal processing methods illustrated in Fig. 4, demonstrate the application variants of various standard radar algorithms under MIMO-OFDM ISAC and analyze their complexity and performance.
@@ IV-A Far-field Sensing Signal Model
For far-field MIMO-OFDM ISAC, the parameters to be estimated for each sensing target are the AoA $\theta_{k}$, propagation delay $\tau_{k}$, and Doppler frequency $\upsilon_{k}$. In the following, we will discuss how various classic radar algorithms for parameter estimation can be applied to far-field MIMO-OFDM ISAC systems, such as the IDFT/DFT and subspace-based methods. IDFT/DFT-based methods like the periodogram [111] can be easily implemented, but the resolution is limited by the array aperture, signal bandwidth and CPI, respectively. On the other hand, subspace-based methods such as MUSIC [79] and ESPRIT [80], can achieve super-resolution parameter estimation, but they usually require high computational complexity. Additionally, CS and sparse recovery methods, such as OMP [81], are also commonly used for parameter estimation. Compared to IDFT/DFT and subspace-based methods, CS achieves higher accuracy under low SNR and sparse channel conditions but incurs higher complexity, especially with large signal dimensions, making it difficult to implement in real time. Moreover, in scenarios with dense targets or clutter, CS methods are prone to locally optimal or suboptimal solutions, hence degrading estimation performance.
@@
Based on the signal model in (21), we below detail the various far-field MIMO-OFDM ISAC signal processing methods illustrated in Fig. 4.
@@ IV-B Decoupled Parameter Estimation
Since the phase variations of the sensing signal in (21) in the angle, delay and Doppler domains are completely decoupled, the three parameters $(\theta_{k},\tau_{k},\upsilon_{k})$ can be estimated separately or sequentially using the parallel and sequential one-domain parameter estimation methods in Fig. 4.
@@ IV-B1 Parallel One-domain Parameter Estimation
AoA estimation problem: In far-field scenarios, the spatial domain processing can only provide AoA information for each target. In this case, the resulting OFDM sensing data in the subcarrier and symbol domains can be treated as snapshots. As such, the $M_{r}\times N\times P$ dimensional signal $\mathbf{Y}_{s}$ in (21) can be re-organized as an $M_{r}\times Q_{N,P}$ dimensional matrix $\mathbf{X}_{\theta}$, with $Q_{N,P}=NP$, where $\mathbf{X}_{\theta}(:,q_{n,p})=\mathbf{Y}_{s}(:,n,p),q_{n,p}=n+pN,n\in[0,N-1],p\in[0,P-1]$. Thus, $\mathbf{X}_{\theta}$ can be equivalently expressed in matrix form as
@@
Delay estimation problem: The random information-bearing symbols $b_{n,p}$ in (21) depend only on subcarrier index $n$ and symbol index $p$, not on antenna index $m$. They affect the manifold structure in the subcarrier and symbol domains, but not in the spatial domain. Thus, unlike AoA estimation in (22), to ensure the manifold structure in the subcarrier and symbol domains, the information-bearing symbols $b_{n,p}$ in (21) need to be removed. The transmitted data is known a priori at the Sen-RX in the monostatic ISAC mode, or it can be demodulated and detected in the bi-static ISAC mode. Thus, the communication symbols can be removed by performing element-wise division on the received signal in (21), i.e.,
@@
where $\mathbf{A}_{\tau}\in\mathbb{C}^{N\times K}$ is the steering matrix in the subcarrier domain given by
@@
and $\mathbf{a}_{\tau}(\tau_{k})\in\mathbb{C}^{N\times 1}$ is the steering vector in the delay domain given by
@@
Besides, the $(k,q_{m,p})$-th element of $\mathbf{S}_{\tau}\in\mathbb{C}^{K\times Q_{M_{r},P}}$ is
@@
The delay estimation problem is to estimate the propagation delays $\mathbb{S}_{\tau}=\{\tau_{1},\cdots,\tau_{K}\}$ of the $K$ sensing targets embedded in steering matrix $\mathbf{A}_{\tau}$, given data matrix $\mathbf{X}_{\tau}$ in (25).
Doppler estimation problem: For Doppler estimation, the data in the antenna and subcarrier domains can be treated as snapshots. The signal after element-wise symbol division in (LABEL:d25) can be restructured into a $P\times Q_{M_{r},N}$ dimensional matrix $\mathbf{X}_{\upsilon}$, with $Q_{M_{r},N}=M_{r}N$, where $\mathbf{X}_{\upsilon}(:,q_{m,n})={\overline{\mathbf{Y}}_{s}(m,n,:)},q_{m,n}=m+nM_{r}$. Then, matrix $\mathbf{X}_{\upsilon}$ can be written as
@@
where $\mathbf{Z}_{\upsilon}$ is corresponding noise, and $\mathbf{A}_{\upsilon}\in\mathbb{C}^{P\times K}$ is the steering matrix in the Doppler domain, given by
@@
and $\mathbf{a}_{\upsilon}(\upsilon_{k})\in\mathbb{C}^{P\times 1}$ is the steering vector in the symbol domain given by
@@
The Doppler estimation problem is to estimate the Doppler frequencies $\mathbb{S}_{\upsilon}=\{\upsilon_{1},\cdots,\upsilon_{K}\}$ of the $K$ sensing targets based on steering matrix $\mathbf{A}_{\upsilon}$, given data matrix $\mathbf{X}_{\upsilon}$ in (29).
@@
There are a total of $K_{\hat{\theta}}K_{\hat{\tau}}K_{\hat{\upsilon}}$ possible parameter groups, obtained as $\left(\hat{\theta}_{k_{\hat{\theta}}},\hat{\tau}_{k_{\hat{\tau}}},\hat{\upsilon}_{k_{\hat{\upsilon}}}\right),\forall k_{\hat{\theta}}=1,\cdots,K_{\hat{\theta}},k_{\hat{\tau}}=1,\cdots,K_{\hat{\tau}},k_{\hat{\upsilon}}=1,\cdots,K_{\hat{\upsilon}}$. In the signal correlation method, the reconstructed signal $\hat{\mathbf{Y}}_{k_{\hat{\theta}},k_{\hat{\tau}},k_{\hat{\upsilon}}}\in\mathbb{C}^{M_{r}\times N\times P}$ for parameter group $\left(\hat{\theta}_{k_{\hat{\theta}}},\hat{\tau}_{k_{\hat{\tau}}},\hat{\upsilon}_{k_{\hat{\upsilon}}}\right)$ is expressed as
@@
The correct parameter groups will exhibit a strong correlation between the reconstructed signal and signal $\overline{\mathbf{Y}}_{s}$. Thus, the $K$ parameter pairs $\left(\hat{\theta}_{k_{\hat{\theta}}},\hat{\tau}_{k_{\hat{\tau}}},\hat{\upsilon}_{k_{\hat{\upsilon}}}\right)$ with the highest correlation coefficient $\hat{C}$ correspond to the $K$ sensing targets.
For power based detection, the power of $\overline{\mathbf{Y}}_{s}$ in (LABEL:d25) in the angle-delay-Doppler domain at the location of each parameter group $\left(\hat{\theta}_{k_{\hat{\theta}}},\hat{\tau}_{k_{\hat{\tau}}},\hat{\upsilon}_{k_{\hat{\upsilon}}}\right)$ can be obtained as
@@
where $\mathbf{a}_{\mathrm{R}}(\hat{\theta}_{k_{\hat{\theta}}})$, $\mathbf{a}_{\tau}(\hat{\tau}_{k_{\hat{\tau}}})$, and $\mathbf{a}_{\upsilon}(\hat{\upsilon}_{k_{\hat{\upsilon}}})$ are the steering vectors in the spatial, delay, and Doppler domains in (18), (27), and (31), respectively. The $K$ parameter groups $\left(\hat{\theta}_{k_{\hat{\theta}}},\hat{\tau}_{k_{\hat{\tau}}},\hat{\upsilon}_{k_{\hat{\upsilon}}}\right)$ with the highest power $\hat{P}$ are taken to be the parameters of the $K$ sensing targets.
@@ IV-B2 Sequential One-domain Parameter Estimation
The sequential one-domain parameter estimation method, as shown in Fig. 4, estimates the target parameters in sequence over the three domains. It typically begins in the domain with the highest resolution so that more targets can be distinguished. For example, consider a case in which the array has many elements, while the bandwidth is narrow and the CPI is short. In this case, the angular resolution is highest, followed by the delay and Doppler resolutions, so it is best to first estimate the AoA of the sensing targets using $\mathbf{X}_{\theta}$ in (22), and then use beamforming to extract the signal of targets at different angles. Then, the delays are estimated for targets at each angle. Since there may be multiple targets at the same angle, beamforming (matched filtering) can be applied in the delay domain to further distinguish targets with different delays. Finally, the corresponding Doppler frequency shift is estimated for targets at each delay. Thus, the target parameters in the three domains are directly grouped in this sequential method.
Specifically, as shown in Fig. 4, assume that a total of $K_{\hat{\theta}}$ AoAs are estimated as $\hat{\theta}_{k_{\hat{\theta}}},k_{\hat{\theta}}=1,\cdots,K_{\hat{\theta}}$, based on (22), where $K_{\hat{\theta}}\leq K$ since not all targets are distinguishable in the spatial domain. Let $N_{k_{\hat{\theta}}}\geq 1$ denote the number of targets at each estimated angle and $K=\sum_{{k_{\hat{\theta}}}=1}^{K_{\hat{\theta}}}N_{k_{\hat{\theta}}}$. Before performing delay domain estimation, we first apply spatial domain beamforming based on the estimated AoAs in order to exploit the beamforming gain and to decouple the signals in the spatial domain. After performing beamforming on (LABEL:d25), the resulting signal $\mathbf{X}_{k_{\hat{\theta}}}\in\mathbb{C}^{N\times P},k_{\hat{\theta}}=1,\cdots,K_{\hat{\theta}}$, is
@@
where $\Omega_{k_{\hat{\theta}}}=\left\{k|\theta_{k}\approx\hat{\theta}_{k_{\hat{\theta}}}\right\}$, $\tilde{\alpha}_{i,{k_{\hat{\theta}}}}=\overline{\alpha}_{i,{k_{\hat{\theta}}}}\mathbf{r}^{H}(\hat{\theta}_{k_{\hat{\theta}}})\mathbf{a}_{\mathrm{R}}(\theta_{i,{k_{\hat{\theta}}}})$ is the resulting complex channel gain, $\tau_{i,{k_{\hat{\theta}}}}$ and $\upsilon_{i,{k_{\hat{\theta}}}}$ are the delay and Doppler of the $i$-th target at the estimated angle $\hat{\theta}_{k_{\hat{\theta}}}$, and $\mathbf{r}\Big{(}\hat{\theta}_{k_{\hat{\theta}}}\Big{)}$ is the beamforming vector used in the spatial domain for the estimated angle $\hat{\theta}_{k_{\hat{\theta}}}$.
There are three classic beamforming methods that are used to achieve a balance between complexity and performance. These methods include maximum-ratio combining (MRC), zero-forcing (ZF), and MMSE. The expressions for the corresponding beamforming vectors are given by
$\mathbf{r}_{\mathrm{MRC}}\Big{(}\hat{\theta}_{{k_{\hat{\theta}}}}\Big{)}$ $\triangleq{\mathbf{a}_{\mathrm{R}}\left(\hat{\theta}_{{k_{\hat{\theta}}}}\right)}\big{/}{\left\|\mathbf{a}_{\mathrm{R}}\left(\hat{\theta}_{{k_{\hat{\theta}}}}\right)\right\|},$ (37)
$\mathbf{r}_{\mathrm{ZF}}\Big{(}\hat{\theta}_{{k_{\hat{\theta}}}}\Big{)}$ $\triangleq{\widehat{\mathbf{A}}_{{k_{\hat{\theta}}}}\mathbf{a}_{\mathrm{R}}\left(\hat{\theta}_{{k_{\hat{\theta}}}}\right)}\big{/}{\left\|\widehat{\mathbf{A}}_{{k_{\hat{\theta}}}}\mathbf{a}_{\mathrm{R}}\left(\hat{\theta}_{{k_{\hat{\theta}}}}\right)\right\|},$
$\mathbf{r}_{\mathrm{MMSE}}\Big{(}\hat{\theta}_{{k_{\hat{\theta}}}}\Big{)}$ $\triangleq{\mathbf{C}_{{k_{\hat{\theta}}}}^{-1}\mathbf{a}_{\mathrm{R}}\left(\hat{\theta}_{{k_{\hat{\theta}}}}\right)}\big{/}{\left\|\mathbf{C}_{{k_{\hat{\theta}}}}^{-1}\mathbf{a}_{\mathrm{R}}\left(\hat{\theta}_{{k_{\hat{\theta}}}}\right)\right\|},$
where $\widehat{\mathbf{A}}_{{k_{\hat{\theta}}}}\triangleq\mathbf{I}_{M_{r}}-\mathbf{A}_{k_{\hat{\theta}}}\left(\mathbf{A}_{k_{\hat{\theta}}}^{H}\mathbf{A}_{k_{\hat{\theta}}}\right)^{-1}\mathbf{A}_{k_{\hat{\theta}}}^{H}$, $\mathbf{A}_{{k_{\hat{\theta}}}}=\left[\mathbf{a}_{\mathrm{R}}\left(\hat{\theta}_{1}\right),\cdots,\mathbf{a}_{\mathrm{R}}\left(\hat{\theta}_{{k_{\hat{\theta}}}-1}\right),\mathbf{a}_{\mathrm{R}}\left(\hat{\theta}_{{k_{\hat{\theta}}}+1}\right),\cdots,\mathbf{a}_{\mathrm{R}}\left(\hat{\theta}_{K_{\hat{\theta}}}\right)\right]$, and $\mathbf{C}_{{k_{\hat{\theta}}}}=\sum_{k\neq{k_{\hat{\theta}}}}^{K_{\hat{\theta}}}\frac{|\overline{\alpha}_{k}|^{2}}{\sigma^{2}}\mathbf{a}_{\mathrm{R}}\left(\hat{\theta}_{k}\right)\mathbf{a}_{\mathrm{R}}^{H}\left(\hat{\theta}_{k}\right)+\mathbf{I}_{M_{r}}$. In the delay and Doppler domains, similar beamformers can be employed.
@@
where $\tilde{\mathbf{\alpha}}=[\tilde{\alpha}_{1},\cdots,\tilde{\alpha}_{N_{k_{\hat{\theta}}}}]^{T}$ is the vector of resulting complex-valued gains, $\mathbf{A}_{\tau,{k_{\hat{\theta}}}}=[\mathbf{a}_{\tau}\begin{pmatrix}\tau_{1}\end{pmatrix},\cdots,\mathbf{a}_{\tau}(\tau_{N_{k_{\hat{\theta}}}})]$ and $\mathbf{A}_{\upsilon,{k_{\hat{\theta}}}}=[\mathbf{a}_{\upsilon}(\upsilon_{1}),\cdots,\mathbf{a}_{\upsilon}(\upsilon_{N_{k_{\hat{\theta}}}})]$ are the steering matrices in the delay and Doppler domains for the targets located at angle $\hat{\theta}_{k_{\hat{\theta}}}$. Then, the matrices $\mathbf{X}_{k_{\hat{\theta}}},{k_{\hat{\theta}}}=1,\cdots,K_{\hat{\theta}}$, can be exploited to estimate the delays $\mathbb{S}_{\tau,{k_{\hat{\theta}}}}=\left\{\tau_{i,k_{\hat{\theta}}}|i=1,\cdots,{N_{k_{\hat{\theta}}}}\right\}$ of the targets at different angles. Furthermore, for targets with different delays at each angle, further target discrimination can be achieved through beamforming in the delay domain. In particular, the signal for a target whose AoA and delay are estimated as $(\hat{\theta}_{k_{\hat{\theta}}},\hat{\tau}_{i,{k_{\hat{\theta}}}})$ can be reduced to $\mathbf{x}_{i,{k_{\hat{\theta}}}}^{T}=\mathbf{r}^{H}\left(\hat{\tau}_{i,{k_{\hat{\theta}}}}\right)\mathbf{X}_{k_{\hat{\theta}}}$, $i=1,\cdots,N_{k_{\hat{\theta}}}$. Then, the Doppler $\hat{\upsilon}_{i,{k_{\hat{\theta}}}}$ can be estimated from ${\mathbf{x}}_{{k_{\hat{\theta}}}}\in\mathbb{C}^{P\times 1}$ for the target.

### v2

@@ IV Far-field MIMO-OFDM ISAC
In this section, we elaborate on how to sense far-field targets based on the signal processing methods illustrated in Fig. 6, demonstrate the application of various standard radar algorithms under MIMO-OFDM ISAC, and analyze their complexity and performance.
@@ IV-A Far-field Sensing Signal Model
For far-field MIMO-OFDM ISAC, the parameters to be estimated for each sensing target are the AoA $\theta_{k}$, propagation delay $\tau_{k}$, and Doppler frequency $\upsilon_{k}$. In the following, we will discuss how various classic radar algorithms for parameter estimation can be applied to far-field MIMO-OFDM ISAC systems, such as the IDFT/DFT and subspace-based methods. IDFT/DFT-based methods like the periodogram [122] can be easily implemented, but the resolution is limited by the array aperture, signal bandwidth, and CPI, respectively. On the other hand, subspace-based methods such as MUSIC [79] and ESPRIT [80] can achieve super resolution but they usually require high computational complexity. Additionally, CS and sparse recovery methods such as OMP [81] are also commonly used for parameter estimation. Compared to IDFT/DFT and subspace-based methods, CS achieves higher accuracy under low SNR and sparse channel conditions but incurs higher complexity, especially with large signal dimensions, making it difficult to implement in real time. Moreover, in scenarios with dense targets or clutter, CS methods are prone to locally optimal or suboptimal solutions, hence degrading estimation performance.
@@
Based on the signal model in (21), we detail below the various far-field MIMO-OFDM ISAC signal processing methods illustrated in Fig. 6.
@@ IV-B Decoupled Parameter Estimation
Since the phase variations of the sensing signal in (21) in the angle, delay, and Doppler domains are completely decoupled, the three parameters $(\theta_{k},\tau_{k},\upsilon_{k})$ can be estimated separately or sequentially using the parallel and sequential one-domain parameter estimation methods outlined in Fig. 6.
@@ IV-B1 Parallel One-domain Parameter Estimation
AoA estimation: In far-field scenarios, the spatial domain processing can only provide AoA information for each target. In this case, the resulting OFDM sensing data in the subcarrier and symbol domains can be treated as snapshots. As such, the $M_{r}\times N\times P$ dimensional signal $\mathbf{Y}_{s}$ in (21) can be re-organized as an $M_{r}\times Q_{N,P}$ dimensional matrix $\mathbf{X}_{\theta}$, with $Q_{N,P}=NP$, where $\mathbf{X}_{\theta}(:,q_{n,p})=\mathbf{Y}_{s}(:,n,p),q_{n,p}=n+pN,n\in[0,N-1],p\in[0,P-1]$. Thus, $\mathbf{X}_{\theta}$ can be equivalently expressed in matrix form as
@@
Delay estimation: The random information-bearing symbols $b_{n,p}$ in (21) depend only on subcarrier index $n$ and symbol index $p$, not on antenna index $m$. They affect the manifold structure in the subcarrier and symbol domains, but not the spatial domain. Thus, unlike AoA estimation in (22), to ensure the manifold structure in the subcarrier and symbol domains, the information-bearing symbols $b_{n,p}$ in (21) need to be removed. The transmitted data is known a priori at the Sen-RX in the monostatic case, or it can be demodulated and detected in the bistatic ISAC mode. The communication symbols can be removed by performing element-wise division on the received signal in (21), i.e.,
@@
where $\mathbf{A}_{\tau}\in\mathbb{C}^{N\times K}$ is the steering matrix in the subcarrier domain
@@
and $\mathbf{a}_{\tau}(\tau_{k})\in\mathbb{C}^{N\times 1}$ is the steering vector in the delay domain
@@
The $(k,q_{m,p})$-th element of $\mathbf{S}_{\tau}\in\mathbb{C}^{K\times Q_{M_{r},P}}$ is
@@
For the model in (25), the goal is to estimate the propagation delays $\mathbb{S}_{\tau}=\{\tau_{1},\cdots,\tau_{K}\}$ of the $K$ sensing targets embedded in steering matrix $\mathbf{A}_{\tau}$, given data matrix $\mathbf{X}_{\tau}$ in (25).
Doppler estimation: For Doppler estimation, the data in the antenna and subcarrier domains can be treated as snapshots. The signal after element-wise symbol division in (LABEL:d25) can be restructured into a $P\times Q_{M_{r},N}$ dimensional matrix $\mathbf{X}_{\upsilon}$, with $Q_{M_{r},N}=M_{r}N$, where $\mathbf{X}_{\upsilon}(:,q_{m,n})={\overline{\mathbf{Y}}_{s}(m,n,:)},q_{m,n}=m+nM_{r}$. Then, matrix $\mathbf{X}_{\upsilon}$ can be written as
@@
where $\mathbf{Z}_{\upsilon}$ is the corresponding noise, $\mathbf{A}_{\upsilon}\in\mathbb{C}^{P\times K}$ is the steering matrix in the Doppler domain
@@
and $\mathbf{a}_{\upsilon}(\upsilon_{k})\in\mathbb{C}^{P\times 1}$ is the steering vector in the symbol domain
@@
For the model in (29), the goal is to estimate the Doppler frequencies $\mathbb{S}_{\upsilon}=\{\upsilon_{1},\cdots,\upsilon_{K}\}$ of the $K$ sensing targets based on steering matrix $\mathbf{A}_{\upsilon}$, given data matrix $\mathbf{X}_{\upsilon}$ in (29).
@@
[…]
@@
[…]
@@
[…]
@@ IV-B2 Sequential One-domain Parameter Estimation
The sequential one-domain parameter estimation method, as outlined in Fig. 6, sequentially estimates the target parameters over the three domains. The approach typically begins in the domain with the highest resolution so that more targets can be distinguished. Consider a case in which the array has many elements, while the bandwidth is narrow and the CPI is short. In this case, the angular resolution is highest, followed by the delay and Doppler resolutions, so it is best to first estimate the AoA of the sensing targets using $\mathbf{X}_{\theta}$ in (22), and then use beamforming to extract the target reflections at different angles. Then, the delays are estimated for targets at each angle. Since there may be multiple targets at approximately the same angle, beamforming (matched filtering) can be applied in the delay domain to further distinguish targets with different delays. Finally, the corresponding Doppler frequency is estimated for targets at each delay. Thus, the target parameters in the three domains are directly grouped in this sequential method.
More specifically, assume that a total of $K_{\hat{\theta}}$ AoAs are estimated as $\hat{\theta}_{k_{\hat{\theta}}},k_{\hat{\theta}}=1,\cdots,K_{\hat{\theta}}$, based on (22), where $K_{\hat{\theta}}\leq K$ since not all targets are distinguishable in the spatial domain. Let $N_{k_{\hat{\theta}}}\geq 1$ denote the number of targets at each estimated angle and $K=\sum_{{k_{\hat{\theta}}}=1}^{K_{\hat{\theta}}}N_{k_{\hat{\theta}}}$. Before performing delay domain estimation, we first apply beamforming based on the estimated AoAs in order to exploit the beamforming gain and decouple the signals in the spatial domain. After applying beamforming to (LABEL:d25), the resulting signal $\mathbf{X}_{k_{\hat{\theta}}}\in\mathbb{C}^{N\times P},k_{\hat{\theta}}=1,\cdots,K_{\hat{\theta}}$, is
@@
[…]
Classical beamforming methods that can be used in this approach include maximum-ratio combining (MRC), zero-forcing (ZF), and minimum mean-square error (MMSE) beamforming. The expressions for the corresponding beamforming vectors are given by
[…]
@@
[…]